# Chemical enrichment of metal-poor stars orbiting massive black hole companions

### v2

@@ Abstract
There are millions of undetected black holes wandering through our galaxy. Observatories like Chandra, LIGO, and more recently, Gaia, have provided valuable insights into the configurations of these elusive objects when residing in binary systems. Motivated by these advances, we study, for the first time, the enhanced accretion of metals from the interstellar medium (ISM) onto low-mass companions in binary systems with highly unequal mass ratios, utilizing a series of hydrodynamical simulations. Our study demonstrates that a stellar companion’s metal accretion history from the ISM alone, from its formation to the present, can significantly influence its surface abundances, especially when enhanced by a massive black hole companion. However, this effect is likely only measurable in stars that are still in the main sequence. Once a stellar companion evolves off the main sequence, similar to what has been observed with the Gaia BH3 companion, the initial dredge-up process are likely to erase any excess surface abundance resulting from the metals that were accreted. As we discover more unequal mass ratio binary systems, it is crucial to understand how the observed metallicity of sun-like companions may differ from their birth metallicity, especially if they are not yet evolved.
@@
software: FLASH 4.3 (B. Fryxell et al., 2000), MESA (J. R. Buchler & W. R. Yueh, 1976; G. M. Fuller et al., 1985; C. A. Iglesias & F. J. Rogers, 1993; T. Oda et al., 1994; D. Saumon et al., 1995; C. A. Iglesias & F. J. Rogers, 1996; N. Itoh et al., 1996; K. Langanke & G. Mart´ınez-Pinedo, 2000; F. X. Timmes & F. D. Swesty, 2000; F. J. Rogers & A. Nayfonov, 2002; A. W. Irwin, 2004; J. W. Ferguson et al., 2005; S. Cassisi et al., 2007; A. I. Chugunov et al., 2007; R. H. Cyburt et al., 2010; A. Y. Potekhin & G. Chabrier, 2010; B. Paxton et al., 2011, 2013, 2015, 2018, 2019; A. S. Jermyn et al., 2023), MESA SDK (R. Townsend, 2021), yt (M. J. Turk et al., 2011)
@@ 1 Introduction
In a universe where multiple-star systems are ubiquitous, it becomes essential to understand how the interactions between the companions shape their properties. For the case of massive stars, it is estimated that most of them reside in binaries (e.g., H. Sana et al., 2012). Recent observational studies additionally estimate that 7070% of stars with masses of $M\approx 10\,M_{\odot}$ are found in multi-body systems (M. Moe et al., 2019). Moreover, there is an inverse correlation between stellar metallicity and binary fraction, implying that a larger fraction of lower metallicity stars are found in multiple stellar systems. It has been shown that solar-like stars with metallicities [Fe/H] lower than 0.5-0.5 have a multiplicity fraction $\approx 2-3$ times larger than that of solar metallicity stars (C. Badenes et al., 2018).
@@
Stars are thought to retain a chemical record of their original environments through their current metal abundances (e.g., A. Frebel & J. E. Norris, 2013, 2015). This characteristic makes low-mass, metal-poor stars (i.e., old stars) in the Galaxy important for studying star formation during the early epochs of the Universe (e.g., P. Macias & E. Ramirez-Ruiz, 2018). These stars provide valuable insights into the conditions that existed during the Galaxy’s formation and may reveal characteristics of the earliest metal-poor stars (J. L. Johnson, 2015), which are believed to be the stellar progenitors of black holes observed by LIGO (e.g., T. Kinugawa et al., 2014).
@@
Efforts to analyze the abundance properties of metal-poor stars depend on the assumption that these stars accurately reflect the metal abundance of the star-forming clouds in which they were born (e.g., A. N. Kolborg et al., 2022, 2023). However, this assumption may be incorrect if stars are able to efficiently accrete metals from the interstellar medium (ISM) as they orbit within the Galaxy’s gravitational potential over their lifetimes (e.g., H. Bondi & F. Hoyle, 1944; R. J. Talbot & M. J. Newman, 1977). Given their significant ages, low-mass, metal-poor stars could interact with a substantial volume of ISM gas, depending on their orbital evolution. Over time, the ISM itself becomes enriched with metals produced by other stars (J. P. Naiman et al., 2018b).
@@
The chemical enrichment of stars through accretion from the ISM has been the subject of study for several decades (R. J. Talbot & M. J. Newman, 1977; C. Alcock & A. Illarionov, 1980; Y. Yoshii, 1981; I. Iben, 1983; A. Frebel et al., 2009; Y. Komiya et al., 2009, 2010; J. L. Johnson & S. Khochfar, 2011; K. Hattori et al., 2014; J. L. Johnson, 2015; Y. Komiya et al., 2015). Notably, using the Eris zoom-in cosmological simulation (J. Guedes et al., 2011; S. Shen et al., 2015), which models the assembly of a Milky Way analog, S. Shen et al. (2017) explore the chemical enrichment of sun-like stars (with shallow convective layers) due to the accretion of metal-enriched gas from the ISM during the Galaxy’s development. They concluded that a metal-poor star traveling through the ISM could accumulate enough material throughout its lifetime to potentially increase its birth metallicity. In extreme cases, accretion onto these relatively pristine stars can reach enrichment levels of [Fe/H] $\lesssim-2$, with the median enrichment level being [Fe/H] $\approx-5$ (S. Shen et al., 2017).
@@
Stellar multiplicity offers a unique opportunity to further investigate how the accretion of metal-enriched gas from the ISM can be enhanced by the combined gravitational pull in multiple star systems, particularly those that host massive black hole companions. Most known black holes are identified by the X-rays they emit while consuming material from their stellar companions (e.g., R. A. Remillard & J. E. McClintock, 2006), known as high-mass X-ray binaries (HMXBs). When black holes are more widely separated from their companions, they produce little to no radiation, making them detectable primarily through the gravitational influence they exert on their companions (J. J. Andrews et al., 2019; T. A. Thompson et al., 2019). Remarkably, Gaia has enabled the detection of three quiescent black holes (K. El-Badry et al., 2023a, b; S. Chakrabarti et al., 2023; Gaia Collaboration et al., 2024).
@@
In our Milky Way, most stellar-origin black holes have a mass of about ten solar masses (e.g., F. Özel et al., 2010), with the previous record holder being the Cyg X-1 black hole, estimated to have a mass of around twenty solar masses (J. C. A. Miller-Jones et al., 2021). The heaviest black hole detected in our Milky Way is Gaia BH3 (Gaia Collaboration et al., 2024), which has a mass of $32.7\pm 0.82\,M_{\odot}$, placing it within the mass range observed by LIGO (R. Abbott et al., 2023a). The companion to this black hole is a $0.8M_{\odot}$ G type star with a metallicity of $2.56\pm 0.11$. Only about $1/10^{4}$ stars in the solar neighborhood have metallicities as low as those of Gaia BH3 (K. El-Badry, 2024). Remarkably, the stellar companion has a metallicity lower than that of 99.99% of stars in the solar neighborhood (K. El-Badry, 2024). The metal-poor companion star in the Gaia BH3 system supports the idea that massive black holes form efficiently in metal-poor environments early in cosmic history. While based on a single system, this observation raises the intriguing possibility that black hole companions may be preferentially found among metal-poor stars—a trend that future discoveries will be able to test.
@@
Previous analytical and numerical studies of binary systems have primarily focused on cases with equal mass ratios (A. Antoni et al., 2019; T. A. F. Comerford et al., 2019; N. Kaaz et al., 2023). In this work, we explore, for the first time, the evolution of binaries with unequal mass ratios to understand how their combined gravitational influence on the surrounding gas can significantly increase its original metallicity. This framework is essential for studying the ecosystems of binary systems and the complexities of their interactions.
@@
In this study, we extend both the analytical and numerical frameworks of classical Bondi-Hoyle-Lyttleton (BHL) accretion to binary systems in gaseous media (H. Bondi & F. Hoyle, 1944; F. Hoyle & R. A. Lyttleton, 1939), specifically focusing on mass ratios where $M_{1}/M_{2}\ll 1$. We conduct three-dimensional hydrodynamic simulations with a uniform-density background wind to investigate how the presence of a more massive binary companion affects the accretion rates experienced by the lighter member. Additionally, we integrate our accretion model with the stellar trajectories from the Eris zoom-in cosmological simulation (S. Shen et al., 2017) to analyze the chemical enrichment of low-mass stars in binaries that have massive black hole companions, such as those discovered by Gaia. This paper is structured as follows. In Section 2 we present the analytical formalism to frame our thinking and in Section 3 we introduce our hydrodynamical framework and computational setup. In Section 4 we present the results of our numerical experiments and in Section 5 we use the case study of Gaia BH3 as an illustration of the effects of accretion on low-mass stars when having massive black hole companions. Section 6 summarizes our findings and the implications that the idealized results have on high-mass-ratio binaries.
@@ 2.1 Analytical Derivation
We set up the problem following the Bondi-Hoyle-Lyttleton (BHL) (H. Bondi & F. Hoyle, 1944; F. Hoyle & R. A. Lyttleton, 1939) formalism, where we have the supersonic motion of a massive gravitational particle in a uniform density medium. The gas around the particle gravitationally focuses behind it, allowing it to accrete a substantial amount of material. We characterize the gas by a density $\rho_{\infty}$ and a Mach number $\mathcal{M}=v_{\infty}/c_{\rm{s},\infty}$ where $c_{\rm{s},\infty}$ is the speed of sound in the medium and $v_{\infty}$ is the velocity of the massive particle with respect to the background medium.
@@
which gives us a length scale for the gravitational influence of the particle on the background gas. By setting this length scale, we can calculate the mass accretion rate as the flux of gas through a circular cross section with radius $R_{a}$, that is,
$\dot{M}=\pi R_{a}^{2}\rho_{\infty}v_{\infty}=4\pi G^{2}M^{2}\rho_{\infty}v_{\infty}^{-3}.$ (2)
For a fixed mass $M$ and velocity $v_{\infty}$, we note that $\dot{M}\propto\rho_{\infty}$. For a full review of the Bondi-Hoyle-Lyttleton accretion formalism we refer the reader to R. Edgar (2004). For our case of interest, we generalize the BHL formalism to encompass a binary system comprised of a black hole of mass $M_{\rm{BH}}$ with a companion star of mass $M_{*}$. A key goal of this paper is to understand how $M_{*}$ is affected by the nearby gravitational potential of the black hole. We take the cases where the separation $a\lesssim R_{a}$ such that the black hole and its companion share the same shock structure. We define a mass ratio between the black hole and companion star as $q=M_{*}/M_{\rm{BH}}$. The accretion radius of the binary becomes,
@@
where $v_{*}$ is the velocity of the companion star with respect to the background gas. There are two components that contribute to $v_{*}$, the center of mass velocity of the system, $v_{\infty}$, and the orbital velocity of the star around the center of mass of the binary, $v_{\rm{orb},*}$. The largest of these two velocities will be the dominant term in Eq. 5. That is, if $v_{\infty}\lesssim\sqrt{GM_{\rm{BH}}/a_{0}}$, the orbital velocity dominates. The two velocities are equivalent at a separation of $a_{0}=0.5R_{a}$. If we start with the assumption that $v_{*}=v_{\infty}$, valid for $0.5R_{a}\lesssim a_{0}\lesssim 1.0R_{a}$, we can calculate the mass accretion onto the companion star, which is the flux of material within one accretion radius $R_{a,*}$, for a star in isolation, as
@@
The only difference in dependency that comes into play is the separation, where $\beta=\frac{1}{2}$ instead of 1. The analytical solution indicates that for separations $a_{0}\lesssim 0.5$, the orbital solution is expected to dominate. This assumes the star is moving through unperturbed gas, that is unaffected by the binary’s presence. Instead, the star is embedded in a dense background that has already been gravitationally focused and shocked by the massive black hole. When we look closely at the relative orbital velocity of the star with respect to the gas, we see that it is always lower than $v_{\infty}$ for significantly lower radii than the analytical solution predicts. We show this result in the Appendix A. We therefore conclude that using the relation in Eq. 10 holds true for the majority of close binaries.
@@ 2.2 The Landscape of Unequal Mass Binaries
In an effort to understand the role that massive companions might have on the accretion history of lighter binary members, we explore here the mass ratio distributions in binaries. While the average mass ratio for solar-type binaries is $q\approx 1/2$ (M. Moe et al., 2019), we expect $q$ to go from equal mass, $q=1.0$, to extreme, with $q\lesssim 10^{-4}$, as in the case of extreme mass ratio inspirals around supermassive black holes (P. Amaro-Seoane et al., 2007). Apart from $q$, the binary separations will also range vastly from compact binaries with separations on the order of tens of $R_{\odot}$ to wide binaries with separations larger than millions of AU. To illustrate the diversity of binaries, we plot a selection of representative examples in Figure 1.
Figure 1 shows an illustrative portion of the observed binaries in our local Universe selected from three catalogs: Gaia DR2 wide binaries (WBs) (F. M. Jiménez-Esteban et al., 2019), Gaia DR3 non-single stars (Gaia Collaboration et al., 2023), HMXBs (L. Bildsten et al., 1997; S. S. Tsygankov et al., 2022) and radio pulsar binary systems (T. Rivinius et al., 2020; G. W. Clark, 2000), which we denoted as PSR. In the right panel of Figure 1, we compare the two key length scales of the problem (i.e., $a$ and $R_{a}$). The $y$-axis shows $a$ while the $x$-axis shows $R_{a}$ as defined in Section 2.1, both in AU. The diagonal lines represents fixed separation to the accretion radius values, such that $a/R_{a}=10,1,0.1,\ \text{and}\ 0.01$. Any binaries with $a/R_{a}\gtrsim 1$ will have two independent accretion radii, while any binaries below the $a/R_{a}=1$ line will share an accretion structure. The purple stars at the top left of the plot are Gaia DR2 wide binaries, which are observationally biased towards systems with large separations. Below them, in pink, is a subset of the Gaia DR3 binaries. These stars were selected by a query similar to that of S. Chakrabarti et al. (2023), with the main difference of lowering the restrictions for m2 and rv_nb_transits to be m2_lower > 1 and rv_nb_transits > 5. In dark (light) blue are HMXBs (PSR), which have all shared accretion structures. The yellow, orange and red points are the Gaia black hole systems (K. El-Badry et al., 2023a, b; S. Chakrabarti et al., 2023; Gaia Collaboration et al., 2024), denoted as BH1, BH2 and BH3, respectively. As we show in Section 5.1 with the use of cosmological simulations, we expect the relative velocity of BH3 with respect to the gas to be significantly smaller than its current velocity and, as a result, BH3 is expected to have had a shared accretion structure for the majority of its enhanced mass accretion history.

### v3

@@ Abstract
There are millions of undetected black holes wandering through our galaxy. Observatories like Chandra, LIGO, and more recently, Gaia, have provided valuable insights into the configurations of these elusive objects when residing in binary systems. Motivated by these advances, we study, for the first time, the enhanced accretion of metals from the interstellar medium (ISM) onto low-mass companions in binary systems with highly unequal mass ratios, utilizing a series of hydrodynamical simulations. Our study demonstrates that a stellar companion’s metal accretion history from the ISM alone, from its formation to the present, can significantly influence its surface abundance, especially when enhanced by a massive black hole companion. However, this effect is likely only measurable in stars that are still in the main sequence. Once a stellar companion evolves off the main sequence, similar to what has been observed with the Gaia BH3 companion, the initial dredge-up process are likely to erase any excess surface abundance resulting from the metals that were accreted. As we discover more unequal mass ratio binary systems, it is crucial to understand how the observed metallicity of sun-like companions may differ from their birth metallicity, especially if they are not yet evolved.
@@
software: FLASH 4.3 (Fryxell2000), MESA (Buchler1976; Fuller1985; Iglesias1993; Oda1994; Saumon1995; Iglesias1996; Itoh1996; Langanke2000; Timmes2000; Rogers2002; Irwin2004; Ferguson2005; Cassisi2007; Chugunov2007; Cyburt2010; Potekhin2010; Paxton2011; Paxton2013; Paxton2015; Paxton2018; Paxton2019; Jermyn2023), MESA SDK (Townsend2021), yt (Turk2011)
@@ 1 Introduction
In a universe where multiple-star systems are ubiquitous, it becomes essential to understand how the interactions between the companions shape their properties. For the case of massive stars, it is estimated that most of them reside in binaries (e.g., Sana2012). Recent observational studies additionally estimate that 7070% of stars with masses of $M\approx 10\,M_{\odot}$ are found in multi-body systems (Moe2019). Moreover, there is an inverse correlation between stellar metallicity and binary fraction, implying that a larger fraction of lower metallicity stars are found in multiple stellar systems. It has been shown that solar-like stars with metallicities [Fe/H] lower than 0.5-0.5 have a multiplicity fraction $\approx 2-3$ times larger than that of solar metallicity stars (2018ApJ...854..147B).
@@
Stars are thought to retain a chemical record of their original environments through their current metal abundances (e.g., Frebel_2013; Frebel_2015). This characteristic makes low-mass, metal-poor stars (i.e., old stars) in the Galaxy important for studying star formation during the early epochs of the Universe (e.g., 2018ApJ...860...89M). These stars provide valuable insights into the conditions that existed during the Galaxy’s formation and may reveal characteristics of the earliest metal-poor stars (Johnson_2015), which are believed to be the stellar progenitors of black holes observed by LIGO (e.g., Kinugawa_2014).
@@
Efforts to analyze the abundance properties of metal-poor stars depend on the assumption that these stars accurately reflect the metal abundance of the star-forming clouds in which they were born (e.g., 2022ApJ...936L..26K; 2023ApJ...949..100K). However, this assumption may be incorrect if stars are able to efficiently accrete metals from the interstellar medium (ISM) as they orbit within the Galaxy’s gravitational potential over their lifetimes (e.g., Bondi1944; Talbot_1977). Given their significant ages, low-mass, metal-poor stars could interact with a substantial volume of ISM gas, depending on their orbital evolution. Over time, the ISM itself becomes enriched with metals produced by other stars (2018MNRAS.477.1206N).
@@
The chemical enrichment of stars through accretion from the ISM has been the subject of study for several decades (Talbot_1977; Alcock_1980; Yoshii_1981; Iben_1983; Frebel_2009; Komiya_2009; Komiya_2010; Johnson_2011; Hattori_2014; Johnson_2015; Komiya_2015). Notably, using the Eris zoom-in cosmological simulation (2011ApJ...742...76G; 2015ApJ...807..115S), which models the assembly of a Milky Way analog, Shen_2017 explore the chemical enrichment of sun-like stars (with shallow convective layers) due to the accretion of metal-enriched gas from the ISM during the Galaxy’s development. They concluded that a metal-poor star traveling through the ISM could accumulate enough material throughout its lifetime to potentially increase its birth metallicity. In extreme cases, accretion onto these relatively pristine stars can reach enrichment levels of [Fe/H] $\lesssim-2$, with the median enrichment level being [Fe/H] $\approx-5$ (Shen_2017).
@@
Stellar multiplicity offers a unique opportunity to further investigate how the accretion of metal-enriched gas from the ISM can be enhanced by the combined gravitational pull in multiple star systems, particularly those that host massive black hole companions. Most known black holes are identified by the X-rays they emit while consuming material from their stellar companions (e.g., 2006ARA&A..44...49R), known as high-mass X-ray binaries (HMXBs). When black holes are more widely separated from their companions, they produce little to no radiation, making them detectable primarily through the gravitational influence they exert on their companions (2019ApJ...886...68A; 2019Sci...366..637T). Remarkably, Gaia has enabled the detection of three quiescent black holes (2023MNRAS.518.1057E; 2023MNRAS.521.4323E; 2023AJ....166....6C; gaia_bh3).
@@
In our Milky Way, most stellar-origin black holes have a mass of about ten solar masses (e.g., 2010ApJ...725.1918O), with the previous record holder being the Cyg X-1 black hole, estimated to have a mass of around twenty solar masses (2021Sci...371.1046M). The heaviest black hole detected in our Milky Way is Gaia BH3 (gaia_bh3), which has a mass of $32.7\pm 0.82\,M_{\odot}$, placing it within the mass range observed by LIGO (2023PhRvX..13a1048A). The companion to this black hole is a $0.8M_{\odot}$ G type star with a metallicity of $2.56\pm 0.11$. Remarkably, the stellar companion has a metallicity lower than that of $99.99\%$ of stars in the solar neighborhood, with only $1/10^{4}$ stars having a lower value (2024OJAp....7E..38E). The metal-poor companion star in the Gaia BH3 system supports the idea that massive black holes form efficiently in metal-poor environments early in cosmic history. While based on a single system, this observation raises the intriguing possibility that black hole companions may be preferentially found among metal-poor stars—a trend that future discoveries will be able to test. Furthermore, stars in this mass regime could belong to the “surviving stars” population, a class low mass population III formed in early hydrogen deuteride-cooling clouds (nishijima2023lowmasspopiiistar).
@@
Previous analytical and numerical studies of binary systems have primarily focused on cases with equal mass ratios (Antoni2019; 2019MNRAS.490.5196C; 2023ApJ...944...44K). In this work, we explore, for the first time, the evolution of binaries with unequal mass ratios to understand how their combined gravitational influence on the surrounding gas can significantly increase its original metallicity. This framework is essential for studying the ecosystems of binary systems and the complexities of their interactions.
@@
In this study, we extend both the analytical and numerical frameworks of classical Bondi-Hoyle-Lyttleton (BHL) accretion to binary systems in gaseous media (Bondi1944; hoyle_lyttleton_1939), specifically focusing on mass ratios where $M_{1}/M_{2}\ll 1$. We conduct three-dimensional hydrodynamic simulations with a uniform-density background wind to investigate how the presence of a more massive binary companion affects the accretion rates experienced by the lighter member. Additionally, we integrate our accretion model with the stellar trajectories from the Eris zoom-in cosmological simulation (Shen_2017) to analyze the chemical enrichment of low-mass stars in binaries that have massive black hole companions, such as those discovered by Gaia. This paper is structured as follows. In Section 2 we present the analytical formalism to frame our thinking and in Section 3 we introduce our hydrodynamical framework and computational setup. In Section 4 we present the results of our numerical experiments and in Section 5 we use the case study of Gaia BH3 as an illustration of the effects of accretion on low-mass stars when having massive black hole companions. Section 6 summarizes our findings and the implications that the idealized results have on high-mass-ratio binaries.
@@ 2.1 Analytical Derivation
We set up the problem following the Bondi-Hoyle-Lyttleton (BHL) (Bondi1944; hoyle_lyttleton_1939) formalism, where we have the supersonic motion of a massive gravitational particle in a uniform density medium. The gas around the particle gravitationally focuses behind it, allowing it to accrete a substantial amount of material. We characterize the gas by a density $\rho_{\infty}$ and a Mach number $\mathcal{M}=v_{\infty}/c_{\rm{s},\infty}$ where $c_{\rm{s},\infty}$ is the speed of sound in the medium and $v_{\infty}$ is the velocity of the massive particle with respect to the background medium.
@@
which gives us a length scale for the gravitational influence of the particle on the background gas. By setting this length scale, we can calculate the Bondi mass accretion rate as the flux of gas through a circular cross section with radius $R_{a}$, that is,
$\dot{M_{\rm B}}=\pi R_{a}^{2}\rho_{\infty}v_{\infty}=4\pi G^{2}M^{2}\rho_{\infty}v_{\infty}^{-3}.$ (2)
For a fixed mass $M$ and velocity $v_{\infty}$, we note that $\dot{M}_{\rm B}\propto\rho_{\infty}$. For a full review of the Bondi-Hoyle-Lyttleton accretion formalism we refer the reader to Edgar_2004. For our case of interest, we generalize the BHL formalism to encompass a binary system comprised of a black hole of mass $M_{\rm{BH}}$ with a companion star of mass $M_{*}$. A key goal of this paper is to understand how $M_{*}$ is affected by the nearby gravitational potential of the black hole. We take the cases where the separation $a\lesssim R_{a}$ such that the black hole and its companion share the same shock structure. We define a mass ratio between the black hole and companion star as $q=M_{*}/M_{\rm{BH}}$. The accretion radius of the binary becomes,
@@
where $v_{*}$ is the velocity of the companion star with respect to the background gas. There are two components that contribute to $v_{*}$, the center of mass velocity of the system, $v_{\infty}$, and the orbital velocity of the star around the center of mass of the binary, $v_{\rm{orb},*}$. The largest of these two velocities will be the dominant term in Eq. 5. That is, if $v_{\infty}\lesssim\sqrt{GM_{\rm{BH}}/a}$, the orbital velocity dominates. The two velocities are equivalent at a separation of $a=0.5R_{a}$. If we start with the assumption that $v_{*}=v_{\infty}$, valid for $0.5R_{a}\lesssim a\lesssim 1.0R_{a}$, we can calculate the mass accretion onto the companion star, which is the flux of material within one accretion radius $R_{a,*}$, for a star in isolation, as
@@
The only difference in dependency that comes into play is the separation, where $\beta=\frac{1}{2}$ instead of 1. The analytical solution indicates that for separations $a\lesssim 0.5$, the orbital solution is expected to dominate. This assumes the star is moving through unperturbed gas, that is unaffected by the binary’s presence. Instead, the star is embedded in a dense background that has already been gravitationally focused and shocked by the massive black hole. When we look closely at the relative orbital velocity of the star with respect to the gas, we see that it is always lower than $v_{\infty}$ for significantly lower radii than the analytical solution predicts. We show this result in the Appendix A. We therefore conclude that using the relation in Eq. 10 holds true for the majority of close binaries.
@@ 2.2 The Landscape of Unequal Mass Binaries
In an effort to understand the role that massive companions might have on the accretion history of lighter binary members, we explore here the mass ratio distributions in binaries. While the average mass ratio for solar-type binaries is $q\approx 1/2$ (Moe2019), we expect $q$ to go from equal mass, $q=1.0$, to extreme, with $q\lesssim 10^{-4}$, as in the case of extreme mass ratio inspirals around supermassive black holes (2007CQGra..24R.113A). Apart from $q$, the binary separations will also range vastly from compact binaries with separations on the order of tens of $R_{\odot}$ to wide binaries with separations larger than millions of AU. To illustrate the diversity of binaries, we plot a selection of representative examples in Figure 1.
Figure 1 shows an illustrative portion of the observed binaries in our local Universe selected from three catalogs: Gaia DR2 wide binaries (WBs) (gaia_wide_binaries), Gaia DR3 non-single stars (2023A&A...674A..34G), HMXBs (Bildsten_1997; Tsygankov_2022) and radio pulsar binary systems (Rivinius_1997; Clark_2000), which we denoted as PSR. In the right panel of Figure 1, we compare the two key length scales of the problem (i.e., $a$ and $R_{a}$). The $y$-axis shows $a$ while the $x$-axis shows $R_{a}$ as defined in Section 2.1, both in AU. The diagonal lines represents fixed separation to the accretion radius values, such that $a/R_{a}=10,1,0.1,\ \text{and}\ 0.01$. Any binaries with $a/R_{a}\gtrsim 1$ will have two independent accretion radii, while any binaries below the $a/R_{a}=1$ line will share an accretion structure. The purple stars at the top left of the plot are Gaia DR2 wide binaries, which are observationally biased towards systems with large separations. Below them, in pink, is a subset of the Gaia DR3 binaries. These stars were selected by a query similar to that of 2023AJ....166....6C, with the main difference of lowering the restrictions for m2 (lower confidence level of the secondary mass) and rv_nb_transits (number of transits used to calculate the radial velocity) to be m2_lower > 1 and rv_nb_transits > 5. In dark (light) blue are HMXBs (PSR), which have all shared accretion structures. The yellow, orange and red points are the Gaia black hole systems (2023MNRAS.518.1057E; 2023MNRAS.521.4323E; 2023AJ....166....6C; gaia_bh3), denoted as BH1, BH2 and BH3, respectively. As we show in Section 5.1 with the use of cosmological simulations, we expect the relative velocity of BH3 with respect to the gas to be significantly smaller than its current velocity and, as a result, BH3 is expected to have had a shared accretion structure for the majority of its enhanced mass accretion history.